Circular No. 6151

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

CH CYGNI C. Rossi, Istituto Astronomico, La Sapienza University, Rome; G.

Muratorio, Observatoire de Marseille; and R. Viotti, Istituto Astrofisica Spaziale, CNR, Frascati; communicate: “Preliminary analysis of optical spectrograms (0.02 nm/pixel at 405–533 nm and 0.004 nm/pixel at 652–659 nm), obtained using the AURELIE system of the Observatoire de Haute Provence 1.52-m telescope during Jan. 11.2–16.2 UT, indicates that the high-velocity event reported by Iijima (IAUC 6146) was already present in January. The Balmer lines display prominent, double-peaked emissions (red peak stronger than the blue) and broad, blueshifted P-Cyg absorptions located −450 to −500 km/s from the central dip. The P-Cyg profile appears strongly variable on short timescales (down to 0.4 day). Dramatic variation has been observed in Hβ. On Jan. 14.8, the line displayed a strong absorption component at −1050 km/s, extending from ∼ −900 to −1400 km/s (all velocities are referred to the central absorption); by Jan. 15.8, this line was no longer visible, and another, deep, broad absorption was visible at −540 km/s, extending to −800 km/s, with a shallow emission centered at ∼ −1400 km/s and extending to −2000 km/s. The He I 447-nm line also displayed a broad emission with variable absorption. On Jan. 16.2, the Hα line was characterized by a double-peaked emission with a blue shoulder from −100 to −400 km/s. A shallow P-Cyg absorption extending to < −1000 km/s may also be present, largely masked by the M star’s spectrum. No variation was noted in the low-excitation emission spectrum.”

V1413 AQUILAE U. Munari, T. Tomov, and B. Yudin, Asiago Astrophysical Observa-

tory; and G. Sostero and R. Passuello, Associazione Friulana di Astronomia e Meteorologia, write: “The eclipsing symbiotic nova AS 338 = V1413 Aql entered the still-ongoing ouburst in 1981 (cf. Munari 1992, A.Ap. 257, 163) and it is currently experiencing a new bright maximum after the last one occurred in late 1993 (cf. IAUC 5884). CCD optical spectra secured in March 1995 with the Asiago Observatory 1.82-m and 1.22-m telescopes are dominated by a blue continuum, with only weak Fe II, He I, and hydrogen Balmer lines in emission. Hydrogen lines show the usual double-peaked profile. There is no spectroscopic signature of mass loss from the outbursting component. Photoelectric photometry obtained on Mar. 12.15 UT with the AFAM 0.45-m telescope gives V = 11.3, B −V = +0.9, U −B = −0.6.”

1995 March 20 Daniel W. E. Green